# Continuously tunable uniaxial strain control of van der Waals heterostructure devices

### v1

@@ III Strain calibration
Figure 3: (a, b) Raman spectra of the bowtie-shaped silicon and graphene at different piezo voltages, acquired at a temperature of 2 K. (c, d) The energy of the Raman peak of silicon and graphene as a function of piezo voltage, indicating a linear relationship. (e) Measured strain as a function of piezo voltage. A negative piezo voltage corresponds to the direction of compressive strain. The strain values were estimated from the position of the Raman peaks in (c, d), or determined by gold SG deposited on silicon and the commercial SG glued on the piezo stack. The zero-strain point was calibrated by measuring the Raman spectrum at a strain-free spot, located far from the central bridge of the Si wafer. The Raman peak is at 525.5 cm-1, corresponding to a piezo voltage of nearly -20 volts. By assuming a gauge factor of 2.17 and appropriate strain transmission factor, the strain values obtained from the SGs match well with those measured by Raman spectroscopy, with all exhibiting a linear dependence on the piezo voltage.
@@
Figure 4: (a) Schematic of a twisted bilayer graphene sample, encapsulated in hBN with graphite gates top and bottom gates. The right panel displays a cartoon illustration of the moiré pattern of tBLG. The moiré wavelength is $\lambda_{\rm m}=a/[2\sin(\theta/2)]$, where $a=0.246$ nm is the lattice constant of graphene, and $\theta$ represents the twist angle, respectively. (b) Optical image of a tBLG device shaped into a Hall bar geometry resting on a bowtie-shaped silicon chip. The long axis of the Hall bar is aligned with the uniaxial strain direction, indicated by the red arrows. A nearby gold strain gauge, with a line width of 2 $\mu$m, is highlighted in the white box. The white scale bar is 100 $\mu$m. (c) Resistivity versus band filling factor at different strains for a tBLG device.Tseng et al. (2022) Inset: Zoom-in view of the resistivity nearby $\nu=-2$. (d) Temperature dependence of resistivity at different strains. (e) The relative resistivity changes as a function of strain at several filling factors, showing linearity with strain.
@@
In most piezoelectric strain cell experiments, the strain level of the sample was estimated by measuring the displacement of piezo stacks using a capacitance strain gauge.Park et al. (2020); Bartlett et al. (2021) The measured displacement was then converted to the strain of the sample by using a strain transmission factor or an effective length determined by the finite element analysis. We performed similar measurements and analysis and present the results in appendix. We have also developed two methods to directly measure the amount of strain induced in the vdW heterostructure as we bias the piezos. We discovered that these two methods provides more accurate and reliable measurement of the strain, which we discuss below.
@@
In the first method, we perform Raman spectroscopy on both the silicon wafer and on a monolayer graphene encapsulated in hBN. We use a laser power of 300 $\mu$W, with integration times of 160 seconds for graphene and 10 seconds for silicon. These Raman spectra allow us to precisely determine the magnitude of uniaxial strain independently in the graphene and silicon using the relationships previously established in Refs. Wang et al., 2019 and Ureña, Olsen, and Raskin, 2013. Figures 3(c-d) show the Raman spectra of silicon substrate and graphene taken at different piezo voltages at a temperature of 2 K. Both peaks systematically blueshift due to the effect of compressive uniaxial strain, as show in Figs. 3(e-f). We convert the wavenumber of the Raman peak to uniaxial strain following the previously established relationships, and plot the measured strain value as a function of piezo voltage in Fig. 3(e). The strain induced in the graphene and silicon both vary linearly as a function of piezo voltage, and overlap almost perfectly. The slightly reduced strain in the graphene compared with the silicon indicates that $\approx 80$% of the strain induced in the silicon wafer is transferred to the graphene.
@@
Although Raman spectroscopy provides the most direct means of calibrating the strain in the vdW heterostructure, it is often not possible to perform in-situ optical measurements. For example, transport measurements performed in a dilution refrigerator are generally incompatible with Raman characterization. Therefore, we have also explored electrical means of calibrating the applied strain. The most common technique is to measure the resistance of a long metal wire as the piezo bias is tuned. We have developed custom strain gauges that we fabricate by evaporating a gold meander nearby the vdW heterostructure, as shown in Fig. 2(c). The measured elastoresistance can be converted to strain by
@@
where $GF$ is the gauge factor of the strain gauge. There are, however, two sources of uncertainty inherent to this technique. First, the strain is measured in the silicon and not directly in the vdW material of interest. We have calibrated the typical strain transmission between the silicon substrate and the vdW heterostructure in Appendix B, as measured directly by the gold meander strain gauges, and find that it is typically $\approx 80$%. This is fully consistent with the value extracted from Raman spectroscopy measurements. Second, the gauge factor of the gold strain gauge can vary between $\approx 2.1-3.3$ in our experiments, based on the precise details of the device fabrication (more details can be found in Appendix B). Nevertheless, the evaporated gold strain gauges provide a reliable rough estimate of the induced strain in the vdW heterostructure, and can be used to confirm that the induced strain in the wafer is linear in the applied piezo voltage. Fig. 3(e) shows an example of the measured strain from the gold strain gauge in the same device in which we separately extracted the strain from Raman spectroscopy. We see that both methods are in good agreement.
@@ IV The effect of strain on transport in a twisted bilayer graphene device
In order to demonstrate the power of our strain technique, we measure transport in a vdW heterostructure device consisting of twisted bilayer graphene device as a function of uniaxial strain. The twist angle of the device is $\theta=1.2^{\circ}$, very near the magic angle at which superconductivity and other correlated states are prominent. The rich correlated and topological physics of magic-angle tBLG has been explored in great detail elsewhere; here, we use it as a model system to test our technique, given that the band structure is expected to be highly sensitive to strain. The device consists of tBLG encapsulated by flakes of hBN, all resting on a flake of graphite (further details on this device can be found in Ref. Tseng et al., 2022). The graphite acts as a back gate, and can change the charge carrier density in the tBLG when a voltage between the graphite and tBLG. Figure 4(a) shows a measurement of the longitudinal resistivity, $\rho$ of the device as a function of the band filling factor, $\nu$ ($\nu=4$ corresponds to fully-filled moiré bands, where the factor of four reflects the spin and valley degeneracy of graphene). In our measurements performed at a temperature of 5 K, we see a resistive state at $\nu=0$ indicative of the charge neutrality point, as well as resistive states at other select integer values of $\nu$ indicative of incipient correlated insulating states. As we apply uniaxial strain by changing the bias on the piezostacks, we see that the resistivity of the device can change by hundreds of ohms, depending on the precise value of $\nu$.
The effect of strain is especially prominent around $\nu=-2$, which corresponds to a developing correlated insulating state driven by a spontaneous isospin flavor polarization in the tBLG. Figure 4(d) shows how $\rho$ evolves as both a function of temperature and strain at a fixed band filling factor of $\nu=-1.99$. Overall, the temperature dependence is characteristic of tBLG devices previously reported. By comparing $\rho(T)$ at different values of applied strain, we see that there is a monotonic decrease of the resistivity at all temperatures up to 20 K as compressive strain is applied. Figure 4(e) shows the elastoresistance, $\Delta\rho$, as a function of compressive strain at various $\nu$ nearby $-2$. In all cases, we see a nearly linear change in resistivity as strain is applied. Although future work is necessary to unravel the physics underlying the strain-tuning of these states, the large linear elastoresistance we see clearly illustrates the ability of strain to manipulate the electronic properties of tBLG.
@@ V Conclusion
In summary, we have reported the development of a new technique for applying continuously tunable strain to vdW heterostructure devices of ultra-high quality and arbitrary complexity. The strain device is based on commercially-available three piezostack design, with custom PCBs added in order to electrically connect to vdW heterostructure devices. We achieved large strain in excess of $0.1$% by appropriately modifying the silicon substrate. We are able to demonstrate efficient strain transfer from the piezostacks to the vdW heterostructure. As shown in Fig. 3(e), the measured strain is linear in the applied piezo voltage over the entire range we studied. This linearity implies that the total strain we can achieve can very likely be increased by further improvements in epoxying the silicon wafer to the sample mount, and by applying larger voltages to the piezostacks. Our technique paves the way for future characterization and control of the correlated and topological physics of a variety of vdW heterostructures. The technique is readily compatible with a range of other characterization tools, including scanning probe microscopes.
@@ acknowledgments
This work was mainly supported by the NSF MRSEC at UW under Grant Number DMR-2308979, the Gordon and Betty Moore Foundation’s EPiQS Initiative, grant GBMF6759 to J.-H.C. and Department of Energy, Basic Energy Science Programs under award DE-SC0023062. The vdW heterostructure device fabrication was partially supported by the Army Research Office under Grant Number W911NF-20-1-0211. J.-H.C. acknowledges support from the David and Lucile Packard Foundation. M.Y. and J.-H.C. also acknowledge support from the State of Washington-funded Clean Energy Institute.
@@
Zhaoyu Liu and Xuetao Ma contributed equally to the article.
@@ Appendix A Strain estimation via finite element analysis
In addition to the direct measurement of the strain in the Si substrate and graphene as discussed in the main text, we also estimate the induced strain by measuring the displacement of the piezo stacks and perform finite element analysis. To do this, we epoxy a commercial strain gauge directly atop one of the piezostacks (Fig. 1(b, d)). Since the gauge factor of the commercial strain gauge is well calibrated, we can directly extract the strain in the piezo stacks, $\epsilon_{\rm piezo}$, and the displacement of the piezo stacks, $L_{\rm piezo}\times\epsilon_{\rm piezo}$, where $L_{\rm piezo}=9~{}$mm is the length of the piezostack. The displacement allows us to estimate the strain in the silicon substrate, $\epsilon_{\rm xx}^{\rm norm}$, by the following relationship:
@@
where $L_{\rm gap}=0.5~{}$mm is the the distance between the titanium sample holder and $\mu$ is a dimensionless strain transmission factor that takes into account the strain relaxation due to the epoxy and within the silicon substrate. The factor of 2 takes into account the displacement of both the center piezo stacks and the outer two piezo stacks.
@@
To estimate the strain transmission factor $\mu$, we performed FEA using ANSYS Academic Research Mechanical software to investigate the strain distribution in the bowtie-shaped silicon chip model, as shown in Fig. 5(a). The Young’s modulus and the Poisson’s ratio of the silicon along the <110> directions of a (100) wafer in the simulation are 169 GPa and 0.064, respectively. Hopcroft, Nix, and Kenny (2010) The narrow bridge of the silicon chip is 0.3 mm in width and 50 $\mu$m in thickness. Using $k=Y\frac{A}{L}$, we estimated the effective spring constants of Si and SiO2 layers on a strip substrate matching the dimensions of the central bridge of the bowtie-shaped chip. The Young’s modulus of the SiO2 layer is 70 GPa, leading to an effective spring constant $k=0.012~{}\frac{N}{\mu m}$ that is significantly lower than that of the silicon substrate, $5.1~{}\frac{N}{\mu m}$. Thus, the SiO2 layer can be safely ignored in our FEA simulation.
@@
In this simulation a 5 $\mu$m displacement was applied to both sides of the titanium plate, which results in a total 2% nominal uniaxial strain. We assume the thickness of the epoxy layer to be 50 $\mu$m. In Fig. 5(b, c), we present both the top and cross-sectional views of the strain distribution. Notably, the strain primarily aligns along the white dashed center-line of the silicon wafer. The highest strain region (colored in red) is located at the center, where the vdW heterostructure was fabricated. As shown in Fig. 5(d), in the simulation the strain at the center of the silicon bridge is 1.04%, resulting in a strain transmission factor $\mu_{\rm FEA}=52\%$. Nevertheless, the strain estimated by this method is a factor of two to three times larger than the strain measured by the Raman spectroscopy (Fig. 3(e)). Therefore, we concluded that the direct measurements by either Raman spectroscopy or sputtered gold strain gauges are far more reliable.
@@
Additionally, we simulate the strain at the center of the bridge as a function of the epoxy thickness, $t$. The results, presented in Fig. 5(f), demonstrate that the applied strain gradually decreases as the epoxy thickness exceeds 100 $\mu$m. However, in practice, achieving a precise thickness of epoxy presents a challenge. We attempt to minimize the thickness by using as little epoxy as possible, and by carefully pressing the wafer down in order to extrude excess epoxy out from underneath.
Appendix B Strain transmission calibration

### v2

@@ III Strain calibration
Figure 3: (a) and (b) Raman spectra of the bowtie-shaped silicon and graphene at different piezo voltages, acquired at a temperature of 2 K. (c) and (d) The energy of the Raman peak of silicon and graphene as a function of piezo voltage, indicating a linear relationship. (e) Measured strain as a function of piezo voltage. The negative piezo voltage corresponds to the direction of compressive strain. The strain values were estimated from the position of the Raman peaks in (c) and (d) or determined by the gold SG deposited on silicon and the commercial SG glued on the piezo stack. The zero-strain point was calibrated by measuring the Raman spectrum at a strain-free spot, located far from the central bridge of the Si wafer. The Raman peak is at 525.5 cm-1, corresponding to a piezo voltage of nearly -20 volts. The gauge factor of the gold SG is obtained from the quadratic fitting in Fig. 6(b). The strain values calculated by the SG match well with those measured by Raman spectroscopy, with all exhibiting a linear dependence on the piezo voltage.
@@
Figure 4: (a) Schematic of a twisted bilayer graphene sample, encapsulated in hBN with graphite gates top and bottom gates. The right panel displays a cartoon illustration of the moiré pattern of tBLG. The moiré wavelength is $\lambda_{\rm m}=a/[2\sin(\theta/2)]$, where $a=0.246$ nm is the lattice constant of graphene and $\theta$ represents the twist angle. (b) Optical image of a tBLG device shaped into a Hall bar geometry resting on a bowtie-shaped silicon chip. The long axis of the Hall bar is aligned with the uniaxial strain direction, indicated by the red arrows. A nearby gold strain gauge, with a line width of 2 $\mu$m, is highlighted in the white box. The white scale bar is 100 $\mu$m. (c) Resistivity vs band filling factor at different strains for a tBLG device.Tseng et al. (2022) Inset: Zoom-in view of the resistivity nearby $\nu=-2$. (d) Temperature dependence of resistivity at different strains. (e) The relative resistivity changes as a function of strain at several filling factors, showing linearity with strain.
@@
In most piezoelectric strain cell experiments, the strain level of the sample was estimated by measuring the displacement of piezo stacks using a capacitance strain gauge.Park et al. (2020); Bartlett et al. (2021) The measured displacement was then converted to the strain of the sample by using a strain transmission factor or an effective length determined by the finite element analysis. We performed similar measurements and analysis and present the results in Appendix A. We have also developed two methods to directly measure the amount of strain induced in the vdW heterostructure as we bias the piezo stacks. We discovered that these two methods provide more accurate and reliable measurement of the strain, which we discuss below.
@@
In the first method, we perform Raman spectroscopy on both the silicon wafer and on a monolayer graphene encapsulated in hBN. We use a laser power of 300 $\mu$W, with integration times of 160 seconds for graphene and 10 seconds for silicon. These Raman spectra allow us to precisely determine the magnitude of uniaxial strain independently of graphene and silicon using the relationships previously established in Refs. Wang et al., 2019 and Ureña, Olsen, and Raskin, 2013. Figures 3(c) and 3(d) show the Raman spectra of silicon substrate and graphene taken at different piezo voltages at a temperature of 2 K. Both peaks systematically blueshift due to the effect of compressive uniaxial strain, as shown in Figs. 3(e) and 3(f). We note that the strain cell applies uniaxial stress, which induces both symmetric ($A_{1g}$ irrep) and antisymmetric ($E_{g}$ irrep, in the case of graphene) components of strain. The linear shift of Raman peak energy is due to the symmetric component of strain. The in-plane Poisson’s ratio along <110> directions of a (100) silicon is only 0.064, indicating a nearly uniaxial strain along the stress direction. This uniaxial strain condition is the same as the Raman measurement performed in Ref. Ureña, Olsen, and Raskin, 2013 (Si) and Wang et al., 2019 (graphene), hence we can reliably use the data in previous work to extract the strain value. We convert the wavenumber of the Raman peak to uniaxial strain following the previously established relationships and plot the measured strain value as a function of piezo voltage in Fig. 3(e). We note that a higher bias up to $300$ V, if allowed by the instrumental setup, can be safely applied to the piezo stack at cryogenic temperatures and would likely induce a larger strain. Hicks et al. (2014a); Park et al. (2020) The strain induced in the graphene and silicon both vary linearly as a function of piezo voltage, and overlap almost perfectly. The slightly reduced strain in graphene compared with the silicon indicates that $\approx 80$% of the strain induced in the silicon wafer is transferred to graphene.
@@
Although Raman spectroscopy provides the most direct means of calibrating the strain in the vdW heterostructure, it is often not possible to perform in situ optical measurements. For example, transport measurements performed in a dilution refrigerator are generally incompatible with Raman characterization. Therefore, we have also explored electrical means of calibrating the applied strain. The most common technique is to measure the resistance of a long metal wire as the piezo bias is tuned. We have developed custom strain gauges that we fabricate by evaporating a gold meander nearby the vdW heterostructure, as shown in Fig. 2(c). The measured elastoresistance can be converted to strain by
@@
where $GF$ is the gauge factor of the strain gauge. There are, however, two sources of uncertainty inherent to this technique. First, the strain is measured in silicon and not directly in the vdW material of interest. Second, the gauge factor of the gold strain gauge can vary between $\approx 2.1$ and $3.3$ in our experiments, based on the precise details of the device fabrication (more details can be found in Appendix B). Nevertheless, we observed a systematic variation of the gauge factor as a function of the gauge resistivity. By fitting the gauge factor vs resistivity with a quadratic function, we estimated the gauge factor of the gold strain gauge evaporated near the same device measured by the Rama spectroscopy. Figure 3(e) shows the measured strain from the gold strain gauge compared with the strain measured by the Raman spectroscopy. We see that both methods are in good agreement. Therefore, the evaporated gold strain gauges provide a reliable rough estimate of the induced strain in the vdW heterostructure and can be used to confirm that the induced strain in the wafer is linear in the applied piezo voltage.
@@ IV The effect of strain on transport in a twisted bilayer graphene device
In order to demonstrate the power of our strain technique, we measure the strain-dependent transport of a vdW heterostructure device consisting of a twisted bilayer graphene device. The twist angle of the device is $\theta=1.2^{\circ}$, very near the magic angle at which superconductivity and other correlated states are prominent. The rich correlated and topological physics of magic-angle tBLG has been explored in great detail elsewhere;Balents et al. (2020); Andrei and MacDonald (2020) here, we use it as a model system to test our technique, given that the band structure is expected to be highly sensitive to strain. The device consists of tBLG encapsulated by flakes of hBN, all resting on a flake of graphite (further details on this device can be found in Ref. Tseng et al., 2022). The graphite acts as a back gate and can change the charge carrier density in the tBLG when a voltage is applied between the graphite and tBLG. Figure 4(c) shows a measurement of the longitudinal resistivity, $\rho$ of the device as a function of the band filling factor, $\nu$ ($\nu=4$ corresponds to fully filled moiré bands, where the factor of four reflects the spin and valley degeneracy of graphene). In our measurements performed at a temperature of 5 K, we see a resistive state at $\nu=0$ indicative of the charge neutrality point, as well as resistive states at other selected integer values of $\nu$ indicative of incipient correlated insulating states. As we apply uniaxial strain by changing the bias on the piezo stacks, we see that the resistivity of the device can change by hundreds of ohms, depending on the precise value of $\nu$.
The effect of strain is especially prominent around $\nu=-2$, which corresponds to a developing correlated insulating state at half filling of the moiré valence band in the tBLG. Figure 4(d) shows how $\rho$ evolves both as a function of temperature and strain at a fixed band filling factor of $\nu=-1.99$. Overall, the temperature dependence is characteristic of tBLG devices previously reported.Cao et al. (2020); Polshyn et al. (2019); Jaoui et al. (2022) By comparing $\rho(T)$ at different values of applied strain, we see that there is a monotonic decrease of the resistivity at all temperatures up to 20 K as compressive strain is applied. Figure 4(e) shows the relative resistivity, $\Delta\rho$, as a function of compressive strain at values of $\nu\approx-2$. In all cases, we see a nearly linear change in resistivity as strain is applied. Although future work is necessary to unravel the physics underlying the strain-tuning of these states, the large linear elastoresistance we see clearly illustrates the ability of strain to manipulate the electronic properties of tBLG.
@@ V Conclusion
In summary, we have reported the development of a new technique for applying continuously tunable strain to vdW heterostructure devices of ultrahigh quality and arbitrary complexity. The strain device is based on a commercially available three piezo stack design, with custom PCBs added in order to electrically connect to vdW heterostructure devices. We achieved a large strain in excess of $-0.15$% by appropriately modifying the silicon substrate. We are able to demonstrate efficient strain transfer from the piezo stacks to the vdW heterostructure. As shown in Fig. 3(e), the measured strain is linear in the applied piezo voltage over the entire range we studied. This linearity implies that the total strain we can achieve can very likely be increased by further improvements in epoxying both the top and bottom surfaces of the silicon wafer to the sample mount, and by applying larger voltages to the piezo stacks up to $300$ V. Our technique paves the way for future characterization and control of the correlated and topological physics of a variety of vdW heterostructures. The technique is readily compatible with a range of other characterization tools, including scanning probe microscopes.
@@ acknowledgments
This work was mainly supported by the NSF MRSEC at UW under Grant Number DMR-2308979, the Gordon and Betty Moore Foundation’s EPiQS Initiative, grant GBMF6759 to J.-H.C. and the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Award Number DE-SC0023062. The vdW heterostructure device fabrication was partially supported by the Army Research Office under Grant Number W911NF-20-1-0211. J.-H.C. acknowledges support from the David and Lucile Packard Foundation. M.Y. and J.-H.C. also acknowledge support from the State of Washington-funded Clean Energy Institute. This work made use of shared fabrication facilities provided by NSF MRSEC 1719797.
@@
Zhaoyu Liu and Xuetao Ma contributed equally to this paper.
@@ Appendix A Strain estimation via finite element analysis
In addition to the direct measurement of the strain in the silicon substrate and graphene as discussed in the main text, we also estimate the induced strain by measuring the displacement of the piezo stacks and perform finite element analysis. To do this, we epoxy a commercial strain gauge directly atop one of the piezo stacks [Fig. 1(b) and (d)]. Since the gauge factor of the commercial strain gauge is well calibrated, we can directly extract the strain in the piezo stacks, $\epsilon_{\rm piezo}$, and the displacement of the piezo stacks, $L_{\rm piezo}\times\epsilon_{\rm piezo}$, where $L_{\rm piezo}=9~{}$mm is the length of the piezo stack. The displacement allows us to estimate the strain in the silicon substrate, $\epsilon_{\rm xx}^{\rm norm}$, by the following relationship:
@@
where $L_{\rm gap}=0.5~{}$mm is the distance between the titanium sample holder and $\mu$ is a dimensionless strain transmission factor that takes into account the strain relaxation due to the epoxy and within the silicon substrate. The factor of 2 takes into account the displacement of both the center piezo stacks and the outer two piezo stacks.
@@
To estimate the strain transmission factor $\mu$, we performed FEA using ANSYS Academic Research Mechanical software to investigate the strain distribution in the bowtie-shaped silicon chip model, as shown in Fig. 5(a). Young’s modulus and Poisson’s ratio of the silicon along the <110> directions of a (100) wafer in the simulation are 169 GPa and 0.064, respectively. Hopcroft, Nix, and Kenny (2010) Young’s modulus of the SiO2 layer is 70 GPa. The narrow bridge of the silicon chip is 0.3 mm in width and 50 $\mu$m in thickness. Using $k=Y\frac{A}{L}$, we estimated the effective spring constants of Si and SiO2 layers on a strip substrate matching the dimensions of the central bridge of the bowtie-shaped chip. The effective spring constant of the silicon substrate is $k=5.1~{}N/\mu m$, which is significantly larger that of the SiO2 layer, $0.012~{}N/\mu m$. Hence, the SiO2 layer can be safely ignored in our FEA simulation.
@@
In this simulation, a 5 $\mu$m displacement was applied to both sides of the titanium plate, which results in a total 2% nominal uniaxial strain. We assume the thickness of the epoxy layer to be 50 $\mu$m. In Fig. 5(b, c), we present both the top and cross-sectional views of the strain distribution. Notably, the strain primarily aligns along the white dashed center-line of the silicon wafer. The highest strain region (colored in red) is located at the center, where the vdW heterostructure was fabricated. As shown in Fig. 5(d), in the simulation the strain at the center of the silicon bridge is 1.04%, resulting in a strain transmission factor $\mu_{\rm FEA}=52\%$. Nevertheless, this simulation does not take into account that the titanium strain cell itself may also deform. Using the effective spring constant of $12~{}N/\mu m$ for the strain cellPark et al. (2020) and $5.1~{}N/\mu m$ for the substrate, we estimate that only 70% of the displacement induced by the piezo actuator is transferred to the silicon substrate. Taking this factor into account, the strain estimated by this method is within a factor of two of the strain measured by the Raman spectroscopy [Fig. 3(e)].
@@
Additionally, we simulate the strain at the center of the bridge as a function of the epoxy thickness, $t$. The results, presented in Fig. 5(f), demonstrate that the applied strain gradually decreases as the epoxy thickness exceeds 100 $\mu$m. However, in practice, achieving a precise thickness of epoxy presents a challenge. We attempt to minimize the thickness by using as little epoxy as possible and by carefully pressing the wafer down in order to extrude excess epoxy out from underneath.
Appendix B Gold strain gauge calibration